OpenStudy (anonymous):

Line segment RS is perpendicular to line segment PQ, and the coordinates are R(4, -5), S(-8, 4), P(0, 6), and Q(-3, y). What is the value of y?

OpenStudy (amistre64):

4,-5 8,-4 ---- 12,-9 slope of -9/12 = -3/4

OpenStudy (amistre64):

0,6 3,-y ---- 3, 6-y slope = (6-y)/3

OpenStudy (amistre64):

(6-y) 4 ----- = ---; y = 2 3 3

OpenStudy (amistre64):

0,6 3,-2 ----- 3,4 ; 4/3 * -3/4 = -1 = perps

OpenStudy (anonymous):

so y is 2 ?

OpenStudy (amistre64):

.... yes

OpenStudy (anonymous):

THANK YOU DARLINGS (: yall are awsome (:

OpenStudy (amistre64):

yw :)